OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form

OpenStudy (anonymous):

y+9=x+3 y=x-6

OpenStudy (anonymous):

y=x-6

OpenStudy (anonymous):

y=x-6